# Training Convolutional Neural Networks with the Forward-Forward Algorithm

### 25 Dec 2023 v2

@@ Abstract
The recent successes in analyzing images with deep neural networks are almost exclusively achieved with Convolutional Neural Networks (CNNs). The training of these CNNs, and in fact of all deep neural network architectures, uses the backpropagation algorithm where the output of the network is compared with the desired result and the difference is then used to tune the weights of the network towards the desired outcome. In a 2022 preprint, Geoffrey Hinton suggested an alternative way of training which passes the desired results together with the images at the input of the network. This so called Forward-Forward (FF) algorithm has up to now only been used in fully connected networks. In this paper, we show how the FF paradigm can be extended to CNNs. Our FF-trained CNN, featuring a novel spatially-extended labeling technique, achieves a classification accuracy of 99.16% on the MNIST hand-written digits data set. We show how different hyperparameters affect the performance of the proposed algorithm and compare the results with CNN trained with the standard backpropagation approach. Furthermore, we demonstrate that Class Activation Maps can be used to investigate which type of features are learnt by the FF algorithm.
Keywords: forward-forward, convolutional neural networks, class activation maps
@@ 1 Introduction
Machine learning using deep neural networks (DNN) continues to transform human live in areas as different as art (DALL-E, stable diffusion), medicine (Alpha-Fold), transport (self-driving cars, any time now), or information retrieval (ChatGPT, Bard). Here, the adjective deep refers to the number of layers of artificial neurons which can go up to the hundreds. Training these networks means to shift the weights connecting the layers from their initial, random values to values which produce the correct predictions at the output layer of the DNN. This is achieved with the help of a loss function that computes the aggregate difference between the predicted output and the accurate results, which must be known for the training examples. The algorithm behind the training is some variant of gradient descent: in each round of training each weight is shifted a bit into the direction minimizing the loss by using the derivative of the loss function with respect to that weight. Taking the derivative of a loss function with respect to a given weight is straightforward for a single, output layer. Training the weights of the earlier layers in DNN requires the gradient computation to be performed iteratively by applying the chain rule (Rumelhart et al., 1986). This process is called backpropagation (BP). Due to its importance, the term BP is also often used loosely to refer to the entire learning algorithm including the gradient descent.
@@
Backpropagation, respectively multi-layer gradient descent, has a number of downsides: first, it requires the storage of intermediate results. Depending on the optimizer, the memory consumption of BP is up to 5 times larger than the requirement for storing the weights alone (Hugging Face Community, 2023). This becomes a problem when training large models on GPU cards with limited memory. Second, under the name neuromorphic computing there is an ongoing search of hardware alternatives to CMOS semiconductors, driven by the desire to decrease power consumption and increase processing speed (Christensen et al., 2022). On these new hardware platforms it is often impossible to implement an analog of BP, raising the need for alternative training algorithm. Finally, evolution has clearly developed learning algorithms for neural networks such as our brain. However, those algorithms seem to be quite (but maybe not completely, see e.g. Lillicrap et al., 2020) different from BP. Given the in general high performance of evolutionary solutions to problems, this raises the question if also deep learning could gain from biologically plausible alternatives to BP.
@@
Due to these limitations, there is an ongoing search for alternative training methods. The most radical approach is to abandon the loss function completely and use a new learning paradigm. Neural networks trained with variants of the locally acting Hebbian learning rule (neurons that fire together wire together) have been shown to be competitive with BP (Journé et al., 2023; Zhou, 2022). Another approach, Equilibrium Propagation (Scellier and Bengio, 2017) is a learning framework for energy based models with symmetric connections between neurons. Layer-wise Feedback Propagation (Weber et al., 2023) removes the need for a gradient computation by replacing the objective of reducing the loss with computing a reward signal from the network output and propagating that signal backward into the net.
In contrast, the most conservative approach is to keep gradient descent, but to replace the required gradients with an estimate computed from the difference in loss of two forward passes with slightly modified weights. While the naive version of this approach, labeled zeroth order optimization, can be expected to be extremely inefficient, modern variants seem to be competitive (Baydin et al., 2022; McCaughan et al., 2023; Malladi et al., 2023; Zhao et al., 2023b).
@@
A third category of algorithms maintains the idea to update the weights using derivatives of some signal involving the difference between the present state of the network and the target state. But to relax the requirement to propagate that signal from the output layer backward towards the input. This can either be done by using exclusively an output-derived error signal for training each intermediate layer(Nøkland, 2016; Flügel et al., 2023), or by training each layer with locally available information gathered in two consecutive forward passes. An examples of the latter is the ”Present the Error to Perturb the Input To modulate Activity technique” (PEPITA), which performs the second forward pass with the sum of the input signal used in the first pass and some random projection of the error signal from that pass (Dellaferrera and Kreiman, 2022; Farinha et al., 2023).
@@
Another example of the use of local information gathered in consecutive runs to update the weights is the Forward-Forward (FF) algorithm proposed by Geoffrey Hinton in December 2022. FF training combines two ideas. First, the weights of a given layer are updated using gradients of a locally defined goodness function, which in the original paper is taken to be the sum of the squares of the activities in that layer. Second, the labels are included in the training data which allows the neurons to learn them together. In order to understand which features of the data vote for a given label, half of the data set is comprised of labels combined with wrong images. For this negative data the weights are changed in order to minimize the goodness. In contrast, for the correctly labeled, positive data the weights are modified to maximize the goodness. Both these objectives can be achieved with a local gradient descent, with no need for BP. The term Forward Forward now refers to having two subsequent training steps, one with positive and one with negative data.
When generalizing this training method to multi-layer networks, it is important to assure that each subsequent layer needs to do more than just measure the length of the activity vector of the previous one. This is achieved by using layer normalization (Ba et al., 2016) which normalizes the activity vector for each sample. This is best summarized by the description in Hinton (2022): the activity vector in the first hidden layer has a length and an orientation. The length is used to define the goodness for that layer and only the orientation is passed to the next layer.
@@
There are two ways of using a FF trained network for inference. First, we can simultaneously train a linear classifier using the activities of the neurons in the different layers as input. Alternatively, we can create multiple copies of the data set under consideration and combine each copy with one of the possible labels. The correct label is then the one with the largest goodness during its forward pass. Note that this approach multiplies the amount of computation required for inference with a factor equal to the number of labels.
@@
Given the repute of the proposer, it is not surprising that the Forward Forward algorithm has inspired a number of groups to suggest modified and adapted versions. Examples include the combination with a generative model (Ororbia and Mali, 2023), multiple Convolutional Blocks (not trained with FF) (Zhao et al., 2023a), or extending FF training to Graph Neural Networks (Paliotta et al., 2023) and Spiking Neural Networks (Ororbia, 2023). In accordance with the neuromorphic computation motivation of moving beyond BP, the FF algorithm has also been used to train optical neural networks (Oguz et al., 2023) and microcontroller units with low computational resources (De Vita et al., 2023). Another line of research tries to improve FF by modifications how the goodness is computed (Lee and Song, 2023; Lorberbom et al., 2023; Gandhi et al., 2023), by understanding the sparsity of activation in FF trained networks (Tosato et al., 2023; Yang, 2023), or by exploring the capabilities of FF for self-supervised learning (Brenig and Timofte, 2023).
There has been less activity in terms of practical applications of the FF algorithm. Particularly, in the classification of real world images it has been noted that FF performs worse than BP (Reyes-Angulo and Paheding, 2023) or has to be combined with BP to achieve satisfying results (Paheding and Reyes-Angulo, 2023). We propose that this lack of applications is due to the absence of a method to train the mainstay of modern image processing, Convolutional Neural Networks (CNN) (Rawat and Wang, 2017), with the FF algorithm. The ideas presented here close this gap.
@@
This paper is structured as following: in Section 2 we will first introduce our spatial-extended labeling technique which is crucial to preserve the label information during convolution. Then we will discuss the details of our implementation of the FF-based learning and inference. The results in Section 3 start with a discussion of the optimal results we obtained on the MNIST hand written digits classification. In Section 3.2 we describe our search for the optimal hyperparameters (using validation data). Section 3.3 shows how Class Activation Maps (CAMs) can be used to get a better understanding of the features learned during training. We close in Section 4 with a short discussion.
@@ 2 Materials and Methods
2.1 Spatially-extended Labelling
@@
Fully connected DNN establish connections between each pixel within an image and with each neuron in the next layer. The size of images is often of the order of mega pixels and typical layers have hundreds of neurons. This results in a number of weights to be trained which far exceeds the information available in typical training data sets. In contrast, convolutional layers use small filter kernels, typically sized in the range of 3 by 3 to 7 by 7 pixels, which are applied to all possible positions of the much larger input picture. In this way, each filter creates a new, processed version of the input image. Because the training algorithm only needs to learn the weights in these filter kernels, it is possible to apply hundreds of these kernels in parallel and still have order of magnitudes less parameters to train than in a fully connected network.
@@
FF training requires the labels to be added to the input images. Hinton (2022) achieves this with an one-hot encoding in which the label information is confined to the first 10 pixels in the upper left region of each image. Figure 1 A and B give an example of this technique. However, convolutional layers will not work with this one-hot encoding because for most of the possible positions of the filter kernel, these labels are not part of the input. For CNN it is imperative for the label information to be spatially present over the entire image, ensuring that it is captured by each possible filter position. Moreover, this spatial labeling needs to be homogeneous; concatenated random patterns, as used by Lee and Song (2023), will also not allow for arbitrary filter positions.
@@
Figure 1: Spatially-extended labels are present in the entire image, while one-hot encoding is confined to the upper-left area. For the FF training we need two data sets which both add labels to the images. The top row describes the creation of the positive data set, where the example image is correctly labeled as seven. The bottom row displays an example of the negative data set where the image is combined with a false label (here a one) which was randomly chosen from the 9 possible ones. Left and right of the dashed line we display the two ways of adding the label. (A) and (B) describe the one-hot encoding used in Hinton (2022): the first ten pixels in the top row of the image, which are usually black/zero, are used as indicators. The column number of the single pixels set to 1 corresponds to the target value. (C) and (D) describe the technique used by us in this paper. Each label corresponds to an image of the same size as the input, but with a characteristic gray value wave. The label is included into the image by pixel-wise addition.
Here, we introduce a spatial-extended labeling approach involving the superposition of the training data set image with a second image of identical dimensions. This additional image consists of a gray-value wave with a fixed frequency, phase, and orientation. Each possible label is associated with a distinct configuration of these three parameters.
@@
As indicated in Figure 1 C and D, both the positive and negative data sets are obtained by harnessing this methodology. Labels are created by starting from an empty 2D Fourier space of the same size than the image. We then set a single 2D Fourier mode to a finite amplitude. The mapping between the labels and the wavelength and orientation of the corresponding mode is a matter of choice; we have tested two different options. Set 1, which is reported in Figure 2, is characterized by a combination of 4 different wave orientations (0°, 45°, 90°, 135°) and three different frequencies. Set 2, as shown in Figure 8, Appendix A, uses only one spatial frequency and obtains the different labels from ten equidistant angular orientations in the range [0°, 180°]. The results obtained by Set 2 are slightly worse, we report them in Appendix A.
After the superposition of the label, the images are normalized to the range 0 to 1. The relative contribution $K$ of the label pattern to the total intensity of the image is a hyperparameter whose influence will be described in Section 3.2.3. Note that we choose our negative labels randomly, not specifically hard based on another forward pass, as it was done in Hinton (2022).
@@
Figure 2: Superposition of an image of the digit 7 with the full first set of waves used for the spatially-extended labelling. Only the image with the label 7 is part of the positive data set, while one of the other nine images is randomly selected to be part of the negative data set.
@@ 2.2 Implementation of the Learning Algorithm
The present study uses a network architecture composed of three consecutive FF-trained convolutional layers. All three layers contain the same number of filter matrices which is one of the hyperparameter we examine. We did not add any max pooling layers because we found that those decrease accuracy as described in Appendix C. All neurons use ReLU as activation function.
All results reported are obtained using the MNIST handwritten digits data set (LeCun and Cortes (2010)). We split the original training data into 50,000 images used for training and 10,000 validation images used in the search for the optimal hyperparameters. All reported validation accuracies are averaged over 5 independent runs with standard deviations typically being 0.01%. The original 10,000 test images were only used for obtaining the final result reported in Section 3.1.
Just as in the FF training of DNN, we evaluate the loss of each convolutional layer by computing the sigmoidal function $\sigma$ of the goodness. The goodness is defined as the sum of squared layer activations $y_{i}$, modified by subtracting a user-provided threshold $\theta$. Following the code provided by Geoffrey Hinton [^3] Code is available at https://www.cs.toronto.edu/hinton/ffcode.zip/, and as confirmed by Gandhi et al. (2023), we choose $\theta$ to be equal to the number of neurons $N$ within that layer.
While computing the loss, we have to account for our different objectives regarding positive and negative data:
$\text{loss}_{\text{\,layer}}=\sigma\left(\sum^{N}_{i=1}\begin{cases}y_{i}^{2}-\theta&\text{if positive data}\\ -y_{i}^{2}+\theta&\text{if negative data}\end{cases}\right).$
We do not induce symmetry in our loss as described in Lee and Song (2023). But we took inspiration from Lorberbom et al. (2023) who found improved collaboration between layers by training them with a cumulative network loss, computed by summing the individual layer losses $\text{loss}_{\text{\,layer}}$. Here, we exclude the loss of the first layer since it yielded better accuracy as shown in Appendix B, This also aligns with Hinton’s exclusion of the first layer during the evaluation phase.
@@
$y_{i,\text{norm}}=y_{i}\bigg{/}{\left(\sqrt{\frac{\sum^{N}_{i=1}y_{i}^{2}}{N}}\right)}.$
This assures that each subsequent layer can only use the pattern not the norm of the matrix formed by the activations of the previous layer. Second, the learning rate is modified halfway through the epochs by employing a linear cooldown:
${lr}_{e}=\frac{2\;lr}{E}\left(1+E-e\right),$
where $E$ represents the total number of epochs and $e$ is the current epoch.
In order to study the contribution of the individual layers we define a layer based loss and accuracy which measure only the capability to discriminate between images of the positive and negative data set. We interpret the output of the sigmoid function as a probability, where values $>0.5$ indicate that the layer recognizes the image as belonging to the positive data set. By comparing with the true assignation (positive or negative) we obtain a discrimination accuracy. And by using the probability to compute a binary cross entropy, we compute a layer-specific discrimination loss.

### 07 Jan 2024 v3

@@ Abstract
The recent successes in analyzing images with deep neural networks are almost exclusively achieved with Convolutional Neural Networks (CNNs). The training of these CNNs, and in fact of all deep neural network architectures, uses the backpropagation algorithm, where the output of the network is compared with the desired result, and the difference is then used to tune the weights of the network towards the desired outcome. In a 2022 preprint, Geoffrey Hinton suggested an alternative way of training which passes the desired results together with the images at the input of the network. This so called Forward-Forward (FF) algorithm has up to now only been used in fully connected networks. In this paper, we show how the FF paradigm can be extended to CNNs. Our FF-trained CNN, featuring a novel spatially-extended labeling technique, achieves a classification accuracy of 99.16% on the MNIST hand-written digits data set. We show how different hyperparameters affect the performance of the proposed algorithm and compare the results with CNNs trained with the standard backpropagation approach. Furthermore, we demonstrate that Class Activation Maps can be used to investigate which type of features are learnt by the FF algorithm.
Keywords: forward-forward, backpropagation, convolutional neural networks, class activation maps, MNIST classification.
@@ 1 Introduction
Machine learning using deep neural networks (DNN) continues to transform human life in areas as different as art (DALL-E, stable diffusion), medicine (Alpha-Fold), transport (self-driving cars, any time now), or information retrieval (ChatGPT, Bard). Here, the adjective deep refers to the number of layers of artificial neurons, which can go up to the hundreds. Training these networks means shifting the weights connecting the layers from their initial, random values to values which produce the correct predictions at the output layer of the DNN. This is achieved with the help of a loss function that computes the aggregate difference between the predicted output and the accurate results, which must be known for the training examples. The algorithm behind the training is some variant of gradient descent: in each round of training each weight is shifted a bit into the direction minimizing the loss by using the derivative of the loss function with respect to that weight. Taking the derivative of a loss function with respect to a given weight is straightforward for a single, output layer. Training the weights of the earlier layers in DNNs requires the gradient computation to be performed iteratively by applying the chain rule (Rumelhart et al., 1986). This process is called backpropagation (BP). Due to its importance, the term BP is also often used loosely to refer to the entire learning algorithm including the gradient descent.
@@
Backpropagation, respectively multi-layer gradient descent, has a number of downsides: first, it requires the storage of intermediate results. Depending on the optimizer, the memory consumption of BP is up to 5 times larger than the requirement for storing the weights alone (Hugging Face Community, 2023). This becomes a problem when training large models on GPU cards with limited memory. Second, under the name neuromorphic computing, there is an ongoing search of hardware alternatives to CMOS semiconductors, driven by the desire to decrease power consumption and increase processing speed (Christensen et al., 2022). On these new hardware platforms, it is often impossible to implement an analog of BP, raising the need for alternative training algorithm. Finally, evolution has clearly developed learning algorithms for neural networks such as our brain. However, those algorithms seem to be quite (but maybe not completely, see e.g. Lillicrap et al., 2020) different from BP. Given the in general high performance of evolutionary solutions to problems, this raises the question if deep learning could also gain from biologically plausible alternatives to BP.
@@
Due to these limitations, there is an ongoing search for alternative training methods. The most radical approach is to abandon the loss function completely and use a new learning paradigm. Neural networks trained with variants of the locally acting Hebbian learning rule (neurons that fire together wire together) have been shown to be competitive with BP (Journé et al., 2023; Zhou, 2022). Another approach, Equilibrium Propagation (Scellier and Bengio, 2017) is a learning framework for energy based models with symmetric connections between neurons. Layer-wise Feedback Propagation (Weber et al., 2023) removes the need for a gradient computation by replacing the objective of reducing the loss with computing a reward signal from the network output, and propagating that signal backward into the net. In contrast, the most conservative approach is to keep gradient descent, but to replace the required gradients with an estimate computed from the difference in loss of two forward passes with slightly modified weights. While the naive version of this approach, labeled zeroth order optimization, can be expected to be extremely inefficient, modern variants seem to be competitive (Baydin et al., 2022; McCaughan et al., 2023; Malladi et al., 2023; zhao:23PINN).
@@
A third category of algorithms maintains the idea to update the weights using derivatives of some signal, which involves the difference between the present state of the network and the target state. But it relaxes the requirement to backpropagate that signal from the output layer towards the earlier layers. This can either be done by using exclusively an output-derived error signal for training each intermediate layer (Nøkland, 2016; Flügel et al., 2023), or by training each layer with locally available information gathered in two consecutive forward passes. An example of the latter is the ”Present the Error to Perturb the Input To modulate Activity technique” (PEPITA), which performs the second forward pass with the sum of the input signal used in the first pass and some random projection of the error signal from that pass (Dellaferrera and Kreiman, 2022; Farinha et al., 2023).
@@
Another example of the use of local information gathered in consecutive runs to update the weights is the Forward-Forward (FF) algorithm (Hinton (2022)) proposed by Geoffrey Hinton in December 2022. FF training combines two ideas. First, the weights of a given layer are updated using gradients of a locally defined goodness function, which in Hinton (2022) is taken to be the sum of the squares of the activities in that layer. Second, the labels are included in the training data, which allows the neurons to learn them together. In order to understand which features of the data vote for a given label, half of the data set is comprised of labels combined with wrong images. For this negative data the weights are changed in order to minimize the goodness. In contrast, for the correctly labeled, positive data the weights are modified to maximize the goodness. Both these objectives can be achieved with a local gradient descent, with no need for BP. The term Forward-Forward now refers to having two subsequent training steps, one with positive and one with negative data.
@@
When generalizing this training method to multi-layer networks, it is important to assure that each subsequent layer needs to do more than just measure the length of the activity vector of the previous one. This is achieved by using layer normalization (Ba et al., 2016) which normalizes the activity vector for each sample. This is best summarized by the description in Hinton (2022): the activity vector in the first hidden layer has a length and an orientation. The length is used to define the goodness for that layer and only the orientation is passed to the next layer. There are two ways of using a FF trained network for inference. First, we can simultaneously train a linear classifier using the activities of the neurons in the different layers as input. Alternatively, we can create multiple copies of the data set under consideration and combine each copy with one of the possible labels. The correct label is then the one with the largest goodness during its forward pass. Note that this approach multiplies the amount of computation required for inference with a factor equal to the number of labels.
@@
Given the repute of the proposer, it is not surprising that the Forward-Forward algorithm has inspired a number of groups to suggest modified and adapted versions. Examples include the combination with a generative model (Ororbia and Mali, 2023), multiple convolutional blocks (not trained with FF) (zhao:23ForwardAlgo), or extending FF training to graph neural networks (Paliotta et al., 2023) and spiking neural networks (Ororbia, 2023). In accordance with the neuromorphic computation motivation of moving beyond BP, the FF algorithm has also been used to train optical neural networks (Oguz et al., 2023) and microcontroller units with low computational resources (De Vita et al., 2023). Another line of research tries to improve FF by modifying how the goodness is computed (Lee and Song, 2023; Lorberbom et al., 2023; Gandhi et al., 2023), by understanding the sparsity of activation in FF trained networks (Tosato et al., 2023; Yang, 2023), or by exploring the capabilities of FF for self-supervised learning (Brenig and Timofte, 2023). There has been less activity in terms of practical applications of the FF algorithm. Particularly, in the classification of real world images it has been noted that FF performs worse than BP (Reyes-Angulo and Paheding, 2023) or has to be combined with BP to achieve satisfying results (Paheding and Reyes-Angulo, 2023). We propose that this lack of applications is due to the absence of a method to train the mainstay of modern image processing, Convolutional Neural Networks (CNN) (Rawat and Wang, 2017), with the FF algorithm. The ideas presented here close this gap.
@@
This paper is structured as following: in Section 2 we will first introduce our spatial-extended labeling technique, which is crucial to preserve the label information during convolution. Then we will discuss the details of our implementation of the FF-based learning and inference. The results in Section 3 start with a discussion of the optimal results we obtained on the MNIST hand written digits classification. In Section 3.2, we describe our search for the optimal hyperparameters (using validation data). Section 3.3 shows how Class Activation Maps (CAMs) can be used to get a better understanding of the features learned during training. We close in Section 4 with a short discussion.
@@ 2 Materials and methods
2.1 Spatially-extended labeling
@@
Fully connected DNNs establish connections between each pixel within an image and with each neuron in the next layer. The size of images is often of the order of mega pixels and typical layers have hundreds of neurons. This results in a number of weights to be trained which far exceeds the information available in typical training data sets. In contrast, convolutional layers use small filter kernels, typically sized in the range of 3 by 3 to 7 by 7 pixels, which are applied to all possible positions of the much larger input picture. In this way, each filter creates a new, processed version of the input image. Because the training algorithm only needs to learn the weights in these filter kernels, it is possible to apply hundreds of these kernels in parallel and still have order of magnitudes less parameters to train than in a fully connected network.
@@
FF training requires the labels to be added to the input images. Hinton (2022) achieves this with an one-hot encoding, in which the label information is confined to the first 10 pixels in the upper left region of each image. Figure 1A and Figure 1B give an example of this technique. However, convolutional layers will not work with this one-hot encoding because for most of the possible positions of the filter kernel, these labels are not part of the input. For CNNs, it is imperative for the label information to be spatially present over the entire image, ensuring that it is captured by each possible filter position. Moreover, this spatial labeling needs to be homogeneous; concatenated random patterns, as used by Lee and Song (2023), will also not allow for arbitrary filter positions.
@@
Figure 1: Spatially-extended labels are present in the entire image, while one-hot encoding is confined to the upper-left area. For the FF training we need two data sets, which both add labels to the images. The top row describes the creation of the positive data set, where the example image is correctly labeled as seven. The bottom row displays an example of the negative data set where the image is combined with a false label (here a one) which was randomly chosen from the 9 possible ones. Left and right of the dashed line we display the two ways of adding the label. (A) and (B) describe the one-hot encoding used in Hinton (2022): the first ten pixels in the top row of the image, which are usually black/zero, are used as indicators. The column number of the single pixels set to 1 corresponds to the target value. (C) and (D) describe the technique used in this paper. Each label corresponds to an image of the same size as the input, but with a characteristic gray value wave. The label is included into the image by pixel-wise addition.
@@
Here, we introduce a spatial-extended labeling approach involving the superposition of the training data set image with a second image of identical dimensions. This additional image consists of a gray-value wave with a fixed frequency, phase, and orientation. Each possible label is associated with a distinct configuration of these three parameters. As indicated in Figure 1C and Figure 1D, both the positive and negative data sets are obtained by harnessing this methodology. Labels are created by starting from an empty 2D Fourier space of the same size than the image. We then set a single 2D Fourier mode to a finite amplitude. The mapping between the labels and the wavelength and orientation of the corresponding mode is a matter of choice; we have tested two different options. Set 1, which is reported in Figure 2, is characterized by a combination of 4 different wave orientations (0°, 45°, 90°, 135°) and three different frequencies. Set 2, as shown in Figure 8, Appendix A, uses only one spatial frequency and obtains the different labels from ten equidistant angular orientations in the range [0°, 180°]. The results obtained by set 2 are slightly worse, which we report in Appendix A. After the label superposition, the images are normalized to the range [0, 1]. The relative contribution $K$ of the label pattern to the total intensity of the image is a hyperparameter, whose influence will be described in Section 3.2.3. Note that we choose our negative labels randomly, not specifically hard based on another forward pass, as it suggested in Hinton (2022).
@@
Figure 2: Superposition of an image of the digit 7 with the full first set of waves used for the spatially-extended labeling. Only the image with the label 7 is part of the positive data set, while one of the other nine images is randomly selected to be part of the negative data set.
@@ 2.2 Implementation of the learning algorithm
The present study uses a network architecture composed of three consecutive FF-trained convolutional layers. All three layers contain the same number of filter matrices, which is one of the hyperparameter we examine. We did not add any max pooling layers because we found that those decrease accuracy as described in Appendix C. Additionally, all neurons use ReLU as activation function. All results reported are obtained using the MNIST handwritten digits data set (LeCun and Cortes (2010)). We split the original training data into 50,000 images used for training and 10,000 validation images used in the search for the optimal hyperparameters. All reported validation accuracy values are averaged over 5 independent runs with standard deviations typically being 0.01%. The original 10,000 test images were only used for obtaining the final result reported in Section 3.1.
Just as in the FF training of DNNs, we evaluate the loss of each convolutional layer by computing the sigmoidal function $\sigma$ of the goodness. The goodness is defined as the sum of squared layer activations $y_{i}$, modified by subtracting a user-provided threshold $\theta$. Following the code provided by Geoffrey Hinton [^3] Code available at https://www.cs.toronto.edu/hinton/ffcode.zip/, and as confirmed by Gandhi et al. (2023), we choose $\theta$ to be equal to the number of neurons $N$ within that layer. While computing the loss, we have to account for our different objectives regarding positive and negative data as:
${loss}_{\,layer}=\sigma\left(\sum_{i=1}^{N}\begin{cases}y_{i}^{2}-\theta&\text{if positive data}\\ -y_{i}^{2}+\theta&\text{if negative data}\end{cases}\right).$
Note that we do not induce symmetry in our loss as described in Lee and Song (2023). But we took inspiration from Lorberbom et al. (2023), who found improved collaboration between layers by training them with a cumulative network loss, which was computed by summing the individual layer losses obtained from ${loss}_{\,layer}$. Here, we exclude the loss of the first layer since it yielded better accuracy, as shown in Appendix B. This also aligns with Hinton’s exclusion of the first layer during the evaluation phase Hinton (2022).
@@
$y_{i,norm}=y_{i}\bigg{/}{\left(\sqrt{\frac{\sum^{N}_{i=1}y_{i}^{2}}{N}}\right)}.$
This assures that each subsequent layer can only use the pattern, not the norm of the matrix formed by the activations of the previous layer. Second, the learning rate $lr$ is modified halfway through the epochs by employing a linear cooldown:
${lr}(e)=\frac{2\;{lr}}{E}\left(1+E-e\right),$
where $E$ represents the total number of epochs and $e$ is the current epoch. In order to study the contribution of the individual layers, we define a layer based loss and accuracy which measures only the capability to discriminate between images of the positive and negative data set. We interpret the output of the Sigmoid function as a probability, where values greater than $0.5$ indicate that the layer recognizes the image as belonging to the positive data set. By comparing with the true assignation (positive or negative), we obtain a discrimination accuracy. Lastly, by using the probability to compute a binary cross entropy, we compute a layer-specific discrimination loss.